with the given domain, find the ranges

a) f(x)=3-2x for -2<x<2 (the < have a _ below it)

b) f(x)=x^2 for -5<x<-2

what does the for means. ive been stuck with this two questions for hours any solutions would be appreciated TT


1
Expert's answer
2021-10-04T13:50:58-0400

The range of a function is the complete set of all possible resulting values of the dependent variable.

a)

"f(-2)=3+4=7"

"f(2)=3-4=-1"

Range: "y\isin(-1,7)"


b)

"f(-5)=25"

"f(-2)=4"

Range: "y\isin(4,25)"